# The dual of Philo’s shortest line segment problem

### v2

@@ 1. Introduction
Let $CDE$ be a triangle with acute angles at vertices $D$ and $E$, such that $|CE|>|CD|$. Let $CK$ be the angle bisector of triangle $CDE$. Drop perpendiculars $KL$ and $KM$ to sides $CD$ and $CE$, respectively. Take points $A$ and $B$ on sides $CD$ and $CE$, respectively, such that $|AC|=|LD|$, $|BC|=|ME|$. Drop perpendicular $AA_{0}$, $BB_{0}$, and $CC_{0}$ to side $DE$. Extend line $AB$ to intersect line $DE$ at point $P$. Then $|AA_{0}|+|AP|>|CD|+|CE|$ (see Figure 9).
@@
Given $A(a,b),\ B(c,d),$ with $a<c$ and $b\geq d>0$ find the minimum of $|C_{1}D_{1}|+|C_{1}E_{1}|$, where $C_{1}(x,y)$ is a point such that $y>b$ and rays $C_{1}A$ and $C_{1}B$ intersect line $y=0$ at points $D_{1}$ and $E_{1}$, respectively.
@@
Remark 1. Note that in general $f(x,y)$ is not a convex function. So, we can not use the well know result in analysis which says that a local minimum of a convex function is also the absolute minimum (see e.g. [10], p. 543). Actually, each of functions
@@
and similarly for $f_{2}$. As some numerical examples show there can be points $(x,y)$ for which Hessian of $f=f_{1}+f_{2}$ is also negative.
Remark 2. We proved that triangle $CDE$ has optimal $|CD|+|CE|$ within all other triangles $C_{1}D_{1}E_{1}$, which share common pair of points $A$ and $B$ on their sides. Triangle $CDE$ has also other interesting properties worth mentioning. First drop perpendiculars $AA_{0}$, $BB_{0}$, and $CC_{0}$ to side $DE$. Extend lines $AB$ and $ML$ to intersect line $DE$ at points $P$ and $Q$, respectively (see Figure 9).
@@
  • $|PE|=|QD|=\frac{|DE|^{2}}{|B_{0}C_{0}|-|A_{0}C_{0}|}$. In particular, this implies that if points $D$, $E$, and $C_{0}$ are fixed, and point $C$ moves vertically along $CC_{0}$, then lines $AB$ and $LM$ pass though fixed points $P$ and $Q$, respectively.
  • $|A_{0}B_{0}|=\frac{[CDE]}{R}+|DE|$, where $[CDE]$ and $R$ are the area and the circumradius, respectively, of $\triangle CDE$.
@@
Remark 3. I leave as an open question to prove that if $CK$ in Theorem 2.1 is a symedian, then $|C_{1}D_{1}|^{2}+|C_{1}E_{1}|^{2}>|CD|^{2}+|CE|^{2}$. As a problem for further exploration it would also be interesting to study these questions for more than two points instead of just two points $A$ and $B$. One can also replace line $DE$ with a plane and ask the same question for more than two points in space.
@@ 3. Conclusion
In the paper we solved a geometric optimization problem inspired by Philo’s line. The obtained solution reveals interesting properties of the angle bisector and its projections on the sides of this triangle. The provided proof uses multivariable optimization and geometrical methods. An open problem about symedians is proposed.

### v7

@@ 1. Introduction
Let $CDE$ be a triangle with acute angles at the vertices $D$ and $E$, such that $|CE|>|CD|$. Let $CK$ be the angle bisector of the triangle $CDE$. Drop the perpendiculars $KL$ and $KM$ to the sides $CD$ and $CE$, respectively. Take the points $A$ and $B$ on the sides $CD$ and $CE$, respectively, such that $|AC|=|LD|$, $|BC|=|ME|$. Drop the perpendicular $AA_{0}$, $BB_{0}$, and $CC_{0}$ to side $DE$. Extend line $AB$ to intersect the line $DE$ at the point $P$. Then $|AA_{0}|+|AP|>|CD|+|CE|$ (see Figure 9).
@@
Given $A(a,b),\ B(c,d),$ with $a<c$ and $b\geq d>0$ find the minimum of $|C_{1}D_{1}|+|C_{1}E_{1}|$, where $C_{1}(x,y)$ is a point such that $y>b$ and the rays $C_{1}A$ and $C_{1}B$ intersect the line $y=0$ at the points $D_{1}$ and $E_{1}$, respectively.
@@
Remark 1. Note that in general $f(x,y)$ is not a convex function (cf. [43]). So, we can not use the well know result in analysis which says that a local minimum of a convex function is also the absolute minimum (see e.g. [13], p. 543). Actually, each of the functions
@@
and similarly for $f_{2}$. There are points $(x,y)$ for which Hessian of $f=f_{1}+f_{2}$ is also negative. For example, for the points of the set $\mathbb{D}$ which are on the line $AB$ we can write $x=\left(\lambda+1\right)a-\lambda c$ and $y=(\lambda+1)b-\lambda d$, where $\lambda>0$, and obtain
$(f)_{xx}(f)_{yy}-(f)_{xy}(f)_{yx}=-\frac{\left(\left(\lambda+1\right)^{2}b+d\lambda^{2}\right)^{2}\left(a-c\right)^{2}}{\left(b-d\right)^{4}\lambda^{4}\left(\left(a-c\right)^{2}+\left(b-d\right)^{2}\right)\left(\lambda+1\right)^{4}}<0.$
Similarly, this Hessian is negative for the points of the set $\mathbb{D}$ which are close to the line $y=b$.
Remark 2. We proved that the triangle $CDE$ has optimal $|CD|+|CE|$ within all the other triangles $C_{1}D_{1}E_{1}$, which share common pair of points $A$ and $B$ on their sides. The triangle $CDE$ has also other interesting properties worth mentioning. First drop the perpendiculars $AA_{0}$, $BB_{0}$, and $CC_{0}$ to the side $DE$ and suppose that $|CD|<|CE|$. Extend the lines $AB$ and $ML$ to intersect the line $DE$ at the points $P$ and $Q$, respectively (see Figure 9).
@@
  • $|PE|=|QD|=\frac{|DE|^{2}}{|B_{0}C_{0}|-|A_{0}C_{0}|}$. In particular, this implies that if the points $D$, $E$, and $C_{0}$ are fixed, and the point $C$ moves vertically along $CC_{0}$, then the lines $AB$ and $LM$ pass though the fixed points $P$ and $Q$, respectively.
  • $|DE|=|A_{0}B_{0}|+\frac{[CDE]}{R}$, where $[CDE]$ and $R$ are the area and the circumradius, respectively, of $\triangle CDE$.
@@
Remark 3. It is possible to generalize Problem 1 by asking to find the minimum of $\left(|C_{1}D_{1}|^{p}+|C_{1}E_{1}|^{p}\right)^{\frac{1}{p}}$ $(p\neq 0)$. If $p\geq 1$, then this expression can be interpreted as $l_{p}$ norm ([23], p. 20). If $p\rightarrow 0$, then
$\left(|C_{1}D_{1}|^{p}+|C_{1}E_{1}|^{p}\right)^{\frac{1}{p}}\rightarrow\sqrt{|C_{1}D_{1}|\cdot|C_{1}E_{1}|}.$
We already solved the case $p=1$. In analogy with [33], [43] we can look at the case $p=\infty$. In this case
$\left(|C_{1}D_{1}|^{p}+|C_{1}E_{1}|^{p}\right)^{\frac{1}{p}}=\max{\left(|C_{1}D_{1}|,|C_{1}E_{1}|\right)}$
and the solution is essentially the same. We will only give the answer and some implications.
Problem 2.
Given $A(a,b),\ B(c,d),$ with $a<c$ and $b\geq d>0$ find the minimum of $\max{\left(|C_{1}D_{1}|,|C_{1}E_{1}|\right)}$, where $C_{1}(x,y)$ is a point such that $y>b$ and the rays $C_{1}A$ and $C_{1}B$ intersect the line $y=0$ at the points $D_{1}$ and $E_{1}$, respectively.
Answer. If $k_{0}:=\left(\frac{b+d}{c-a}\right)^{\frac{1}{3}}>\frac{b-d}{c-a}$, then the minimum of function $f(x,y)$ is $f(x^{*},y^{*})$, where
$x^{*}=\frac{a+c}{2}-\frac{b-d}{2k_{0}},\ y^{*}=\frac{b+d}{2}-k_{0}\frac{a-c}{2}.$
If $k_{0}\leq\frac{b-d}{c-a}$, then the minimum of function $f(x,y)$ does not exist. In this case there is infimum of function $f(x,y)$ at the points infinitely close to $(a,b)$ and the value of the infimum is equal to
$b\sqrt{\left(\frac{c-a}{b-d}\right)^{2}+1}.$
Using this result we can prove the following theorem.
Theorem 2.3.
Let $A$ and $B$ be points on the sides $CD$ and $CE$, respectively, of the isosceles triangle $CDE$ such that $\tan{\angle D}=\tan{\angle E}=\left(\frac{|AA_{0}|+|BB_{0}|}{|A_{0}B_{0}|}\right)^{\frac{1}{3}}$, where $AA_{0}$ and $BB_{0}$ are the perpendiculars from points $A$ and $B$, respectively, to the side $DE$, and $|AA_{0}|>|BB_{0}|$ (see Figure 10). Then for any point $C_{1}$ different from $C$, such that the rays $C_{1}A$ and $C_{1}B$ intersect the line $DE$ at the points $D_{1}$ and $E_{1}$, respectively, the inequality $\max{\left(|C_{1}D_{1}|,|C_{1}E_{1}|\right)}>|CD|$ holds true. In particular, if the lines $AB$ and $DE$ intersect at the point $P$, then $|AP|>|CD|$ (cf. [4]).
Figure 10. Inequalities $\max{\left(|C_{1}D_{1}|,|C_{1}E_{1}|\right)}>|CD|$ and $|AP|>|CD|$.
I leave as an open question to study the cases $p=0$, $p=2$ and the other values of $p$. It seems that $p=1$ is a unique case when the inequality is always true.
Problem 3.
(Open) Let $p\neq 1$ be a real number. Let $CK$ be a cevian of the triangle $CDE$ such that $\frac{|DK|}{|KE|}=\left(\frac{|DC|}{|CE|}\right)^{p}$. Drop the perpendiculars $KL$ and $KM$ to the lines $CD$ and $CE$, respectively. Take the points $A$ and $B$ on the lines $CD$ and $CE$, respectively, such that $\overrightarrow{AC}=\overrightarrow{DL}$, $\overrightarrow{BC}=\overrightarrow{EM}$. Take any point $C_{1}$ different from $C$, such that the rays $C_{1}A$ and $C_{1}B$ intersect the line $DE$ at the points $D_{1}$ and $E_{1}$, respectively. Does the inequality
$\left(|C_{1}D_{1}|^{p}+|C_{1}E_{1}|^{p}\right)^{\frac{1}{p}}>\left(|CD|^{p}+|CE|^{p}\right)^{\frac{1}{p}}$
hold true for all possible triangles $CDE$ and points $C_{1}$ (see Figure 2)?
In particular, for the case $p=2$ one can show that if $CK$ of Problem 3 is a symedian, then
$|C_{1}D_{1}|^{2}+|C_{1}E_{1}|^{2}>|CD|^{2}+|CE|^{2}$
is not always true. One can find counterexamples by considering triangles $CDE$ with small $|CD|$ and the point $C_{1}$ close to the point $A$. The case $p<1$ seems to be similar because the point $C$ in this case is not always an absolute extremum point. The expression $\left(|C_{1}D_{1}|^{p}+|C_{1}E_{1}|^{p}\right)^{\frac{1}{p}}$ can get smaller (if $10\leq p<$ 1) and greater (if $p<0$) values when point $C_{1}$ approaches points $A$ or $B$. For example, in the case $p=0$, if $CK$ of Problem 3 is a median, then the inequality $|C_{1}D_{1}|\cdot|C_{1}E_{1}|>|CD|\cdot|CE|$ is not always true. As a problem for further exploration it would also be interesting to study these questions for more than two points instead of just the points $A$ and $B$. One can also replace the line $DE$ with a plane and ask the same question for more than two points in space.
@@ 3. Conclusion
In the paper we solved geometric optimization problems inspired by Philo’s line. The obtained solution reveals interesting properties of the angle bisector and its projections on the sides of this triangle. The provided proof uses multivariable optimization and geometrical methods. Some generalizations and open problems, in particular, the impossibility of the analogous inequality for medians and symedians are discussed.
Appendix
In this part we will give an elementary solution for the classical problem mentioned in the introduction above. In the literature, this problem is usually solved using calculus. A different elementary proof based on angle chasing is given in [8].
Problem 4.
Let the points $D$ and $G$ on the side $EF$ of $\triangle BEF$ such that the points $D$ and $G$ are symmetric with respect to the center of the side $EF$ and $BG\perp EF$. Let $E_{1}F_{1}$ be any line through point $D$ and different from $EF$ such that the points $E_{1}$ and $F_{1}$ are on the rays $BE$ and $BF$, respectively. Prove that $|E_{1}F_{1}|>|EF|$.
Figure 11. Proof of the inequality $|E_{1}F_{1}|>|EF|$.
Solution. Let the lines $BG$ and $E_{1}F_{1}$ meet at the point $S$ (see Figure 11). Let the lines through $E$ and $F$ parallel to $BG$ intersect the line $E_{1}F_{1}$ at the points $T$ and $U$, respectively. Let these parallel lines also intersect the line through $B$ and parallel to $E_{1}F_{1}$ at the points $L$ and $M$, respectively. Then $\frac{|LB|}{|BM|}=\frac{|EG|}{|GF|}=\frac{|DF|}{|DE|}$. By the similarity of $\triangle EE_{1}T$ and $EBL$, $\frac{|E_{1}T|}{|BL|}=\frac{|ET|}{|EL|}$. Analogously, by the similarity of $\triangle FF_{1}U$ and $FBM$, $\frac{|BM|}{|F_{1}U|}=\frac{|MF|}{|UF|}$. Using the similarity of $\triangle ETD$, $\triangle GSD$ and $\triangle FUD$, we find $|UF|=|ET|\cdot\frac{|DF|}{|DE|}$ and $|SG|=|ET|\cdot\frac{|DG|}{|DE|}$. Since $MUSB$ and $SBLT$ are parallelograms, $|MU|=|BS|=|LT|=|BG|-|ET|\cdot\frac{|DG|}{|DT|}$. So, $|MF|=|MU|+|UF|=|BG|+|ET|$ and $|LE|=|LT|-|ET|=|BG|-|ET|\cdot\frac{|EG|}{|DE|}$. By mutiplying the expressions for $\frac{|E_{1}T|}{|BL|}$, $\frac{|LB|}{|BM|}$, and $\frac{|BM|}{|F_{1}U|}$, we obtain $\frac{|E_{1}T|}{|F_{1}U|}=\frac{|BG|+|ET|}{|BG|-|ET|\cdot\frac{|EG|}{|DE|}}>1$. Therefore ${|E_{1}T|}>{|F_{1}U|}$ and consequently, ${|E_{1}F_{1}|}=|E_{1}T|+|TU|-|UF_{1}|>|E_{1}T|+|EF|-|UF_{1}|>|EF|.$ All other positions of the line $E_{1}F_{1}$ and the points $D$ and $G$ can be considered similarly.
@@ Acknowledgments
I thank Pyrkov V.E. for informing me about the existence of the paper [25]. I also thank Sinkevich G.I. for the opportunity to give a talk about the topic in The Seminar on the History of Mathematics [5].